(4pts) Sam is conducting a t test to see if the average study time of all NMSU students is more than 6 hours. 25 students are randomly selected, and the data is collected. After analyzing the data, the test statistic t = 1.78. If significance level is set to be 0.02, which of the following is correct?

A. Reject Ho, but have a chance to commit a type I error.

B. Reject Ho, but have a chance to commit a type II error.

C. Fail to reject Ho, but have a chance to commit a type I error.

D. Fail to reject Ho, but have a chance to commit a type II error.
Answer #1

Here t-critical value at alpha = 0.02 with df = n-1 = 25 - 1 = 2

For one tailed t test is, t-critical = 2.1715

Here t < t-critical

Therefore, we fail to reject H0.

Option (D) is correct.